OpenStudy (anonymous):

Explain how to convert 96 oz to pounds.

OpenStudy (anonymous):

1 lb = 16oz. how many times does 16 divide into 96?

OpenStudy (anonymous):

or use proportion 1/16=x/96 x=96/16 lb

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?